OpenStudy (anonymous):

3x+2y=2 What are possible numbers that can represent x and y?? ~~~URGENT!!!~~~

OpenStudy (anonymous):

please someone...

OpenStudy (y1066380):

x=0,y=1 just random guess

OpenStudy (y1066380):

how many set of numbers do you want?

OpenStudy (anonymous):

x = -2, y = 4

OpenStudy (anonymous):

x = -4, y = 7

OpenStudy (anonymous):

x = -6, y = 10

OpenStudy (anonymous):

I got x=2 and y=-2

OpenStudy (anonymous):

This can keep going on, x decreases by -2, and y increasing by 3.

OpenStudy (lyrae):

The diophantine equation 3x+2y=2 have the following soulutions: \[\large x = 2n, \space y = 1 - 3n, \space n \in \mathbb{Z}\]

OpenStudy (anonymous):

thanks guys, you were all right :)

OpenStudy (anonymous):

x=0, y=1 x=-2, y=4 x=-4, y=7